OpenStudy (anonymous):

Express the area of the region under the curve y = 6x^3 + 7x^2 and above the x-axis in the form B (Integral) f(x) dx A

OpenStudy (anonymous):

I missed this class so I'm lost on how to even approach it :/

OpenStudy (anonymous):

This is a nice question, I like it. Okay to get where it is above the x-axis, you would need to see where it touches the x-axis. Did you graph the function, visualization really helps in this problem.

OpenStudy (anonymous):

I'll graph to right now!

OpenStudy (anonymous):

Okay the points that touch the x-axis is when y = 0, so here we would let y= 0 and the x-values are actually the values that lies on the axis.

OpenStudy (anonymous):

the zero of the function is x=-1.166667

OpenStudy (anonymous):

And 0

OpenStudy (anonymous):

Yes, that is correct. What you got is right. So now if you look at the region under the cure but above the x-axis, it would be from -1.66667 to 0. So in the case b = -1.66667 and a = 0. and f(x) = function given. Hope this helps.

OpenStudy (anonymous):

ohhh okay thank you so much!

OpenStudy (anonymous):

i see to have gotten it wrong ;/

OpenStudy (anonymous):

seem*

OpenStudy (anonymous):

\[\int_{-\frac{7}{6}}^0 \left(6 x^3+7 x^2\right) \, dx=\frac{2401}{2592}=0.93\]